A wet miserable morning at Grosmont today but it did get out much brighter later on in the day and was quite warm and sunny. In attendance were Paul Hutchinson, Angie Buxton, Chris Henwood, Ian Pearson and myself, joined by Jon Bradley and Nigel Hall after lunch.
We arrived to learn that the T2 had been failed the previous evening with leaking small tubes in the firebox. After a cup of tea in the Support Coach and a discussion of the extent and possible cause of the problem, Mark O'Brien went in the firebox to do a visual inspection now the locomotive had cooled down, and reported that nearly all the small tubes below the flue tubes were leaking. Further discussion and inspection during the morning did not cast any more light on the possible cause, but by the end of the day, it had been decided to remove the brick arch, expand and then bead over the tubes, steam test, and then rebuild the brick arch. So probably out of action for a week at least. Discussions will doubtless continue with the MPD as to the possible cause, bearing in mind that the T2 tubes had been tight and given no cause for concern all the way through the operating season to date.
In the meantime, Ian and Chris H set about jobs on the T2 with Chris dismantling the regulator, removing the follower and some of the old gland packing, replacing it with new graphite packing and refitting. Ian took off the steam heat valve which was blowing by, and it also had a blowing joint where it fits to the clack box in the cab. This was taken to the workshop, stripped down and cleaned. The seat was recut and ground in and a new joint made, but the head requires skimming which Paul will do on Saturday and then the valve can be refitted. Angie made and fitted a new piece of wood which supports a bracket securing the low level water filler on the left hand side of the K1 tender. Paul attended to the K1 left hand piston packing which had been giving some trouble recently: it was stripped down, cleaned, adjusted and re- fitted. Back at Deviation Shed, I rubbed down and painted the length of pipe which has been made to fit the bottom of the gate at the North end of No 7 Road to stop small children getting under the gate and falling into the big dark pit. This pipe will be refitted on Saturday. I also did the monthly check of the defibrillator.
Lunch was hosted in the Support Coach today - a bit posher than we are used to! After lunch, Angie removed all the ash from the K1 smoke box and barrowed it to the ash tip. I mixed some cement and grouted around the gate post on No 7 Road to tidy up the area. I also made the afternoon tea. Nigel Hall arrived to collect 2 stand lamps to take to Hopetown, and also helped Jon Bradley clean around the stove area ready for our winter warming.
No work was done on the J27 as it is waiting for a washout and is currently parked outside Deviation Shed on No 8 Road. It is unlikely to work again before the end of the operating season, but will then be in use for the Northern Lights train and the Santas between Pickering and Levisham at the end of the year. With the S160 gear being moved from No 6 Road to the back of No 8 Road now that Hartland has been moved into the fitting shop where it was being prepared for repainting (among other tasks), Chris Cubitt was hoping to move No 29 into Deviation Shed on No 6 Road today for temporary storage. It has a broken spring on the rear bogie and, with no spares available, will have to wait for supplies from the spring manufacturer before it can work again. So, regrettably, there can be no prospect of it double heading the diner with the K1 as a mark of respect to the late John Richardson who was one of the original consortium involved in its purchase. However, the roller shutter door would not wind up as the gears seem to have jammed, and will have to be repaired by a specialist contractor in approximately a week's time. That also saved us having to move the velocipede to give the necessary length for No 29 to fit. We also looked at the piles of material stacked up against the outside of the back wall of the building, and identified a number of items that are clearly scrap and need to be removed to avoid possible damage to the building fabric. The pit pump also seems to need attention as it is very slow at removing water. A good clean out is probably required, a job for which I was volunteered next week!
There were a number of visitors around today, in spite of the weather and time of the year. Locomotives in traffic were Black Five No 5428 Eric Treacy (but sounding a bit off beat), S15 No 825, B1 No 1264 and the DMU.
There will be a small working party on Saturday 19 October 2019 with Paul, Angie and possibly Ian - others will be very welcome to join them though. The next regular working party will be on Wednesday 23 October 2019, but it will not be long now before we start on winter maintenance and return to regular Saturday and Wednesday working parties.